Scot and Vidia, married taxpayers, earn $266,000 in taxable income and $8,900 in interest from an investment in the City of Tampa bonds. Using the U.S. tax rate schedule for married filing jointly, how much federal tax will they owe? What is their average tax rate? What is their effective tax rate? What is their current marginal tax rate? (Do not round intermediate calculations. Round "Average tax rate" & "Effective tax rate" to 2 decimal places)

*Using 2019 Tax Rate Schedules*
Answer #1

As per 2019 tax bracket (married filing jointly):

Federal tax = 19,400 × 10% + (78,950 – 19400 =) 59,550 × 12% + (168,400 – 78,950 =) 89,450 × 22% + (266,000 – 168,400 =) 97,600 × 24%

                        = 1940 + 7146 + 19,679 + 23,424

                        = 52,189 (Answer)

Average tax rate = (Federal tax / Taxable income) × 100

                           = (52,189 / 266,000) × 100

                           = 19.62% (Answer)

Effective tax rate = (Federal tax / Total income) × 100

                             = (52,189 / (266,000 + 8,900)) × 100

                             = (52,189 / 274,900) × 100

                             = 18.98% (Answer)

Marginal tax rate = tax on the last dollar of taxable income

                             = 24% (Answer)